# Provable FDR Control for Deep Feature Selection: Arbitrarily Deep MLPs and Beyond

### v1

@@ Abstract.
We develop a flexible feature selection framework based on deep neural networks that approximately controls the false discovery rate (FDR), a measure of Type-I error. The method applies to architectures whose first layer is fully connected. From the second layer onward, it accommodates multilayer perceptrons of arbitrary width and depth, convolutional and recurrent networks, attention mechanisms, residual connections, and dropout. The procedure also accommodates stochastic gradient descent with data-independent initializations and learning rates. To the best of our knowledge, this is the first work to provide a theoretical guarantee of FDR control for feature selection within such a general deep learning setting.
@@ 1. Introduction
On the statistics side, a central objective has been to provide theoretical guarantees on false discovery rate (FDR) control [abramovich2006adapting, meinshausen2006high, barber2015controlling, candes2018panning, xing2023controlling, dai2023false, dai2023scale, du2023false], often by debiasing sparse estimators like the LASSO [tibshirani1996regression]. These guarantees, albeit rigorous, typically come at the expense of strong modeling assumptions (e.g., the true model follows a generalized linear model, or the feature distribution is known), which can diverge from complex real‑world phenomena and thereby cast doubt on the reliability of the selected features.
On the XAI side, a wealth of attribution methods, e.g., LIME [ribeiro2016should], SHAP [lundberg2017unified], random forest feature importance [breiman2001random], and saliency maps [simonyan2013deep], quantify the contribution of individual features in black‑box models. Thresholding such scores yields a practical selection heuristic, yet without guarantees on Type‑I error; thus, error control has remained elusive.
@@ 1.1. Related works
There are two main research streams of FDR control in feature selection: estimator-based FDR control and knockoff filters. In the former, e.g., for linear regression, whether coefficients are zero or nonzero determines which features should be selected, and FDR control can be achieved by invoking asymptotic normality of the estimators. Representative examples include the Gaussian mirror [xing2023controlling] and data splitting [dai2023false]. A key advantage of these approaches is scale-free property, since they avoid estimating the asymptotic variance, and they have been extended to generalized linear models [dai2023scale]. Similar techniques have also been applied to sliced inverse regression under multi-index models [zhao2022testing], although the theoretical assumptions on the feature dimension and the sparsity level are rather stringent.
As the second line of research, the model-X knockoff [candes2018panning] is groundbreaking in that it can control the FDR without assuming the structure of $\bm{y}\mid\bm{X}$ where $\bm{y}\in\mathbb{R}^{m}$ is a response vector. However, it requires knowledge of the joint distribution of $\bm{X}$, which is restrictive. To address this limitation, methods that estimate the distribution of $\bm{X}$ using generative networks have been proposed [jordon2019knockoffgan, romano2020deep]. In addition, a sequential knockoff sampler for a given feature distribution has been proposed [bates2021metropolized].
@@
Several studies have proposed FDR control methods that employ neural networks, although without theoretical guarantees and typically within restricted classes of architectures. The Neural Gaussian Mirror [xin2020ngm] defines a kernel-based conditional dependence measure and performs feature selection with MLPs. DeepPINK [lu2018deeppink] is a knockoff framework with a specially designed network architecture in which the features are assumed to be jointly Gaussian, and DeepLINK [zhu2021deeplink] relaxes this distributional restriction. These approaches have been shown empirically to achieve FDR control. Nevertheless, because theoretical guarantees are not provided, it remains unclear under what conditions FDR control is achievable, and the range of supported network architectures is limited.
@@ 2. Problem formulation
Suppose that we observe a response vector $\bm{y}=(y_{1},\ldots,y_{m})^{\top}\in\mathbb{R}^{m}$ together with a design matrix $\bm{X}=(\bm{x}_{1},\ldots,\bm{x}_{m})^{\top}\in\mathbb{R}^{m\times n}$, where $m$ is the sample size and $n$ is the number of features. Our goal is to select a relevant subset of feature indices from $[n]$ that are associated with $\bm{y}$. A desirable feature selection procedure controls the false discovery rate (FDR) [benjamini1995controlling], defined by
@@ 3. Theoretical background
The first layer of the network is taken to be dense and fully connected, and we denote its weight matrix by $\bm{W}_{1}\in\mathbb{R}^{n\times q}$. The dependence of the entire network $f_{\mathcal{W}}(\bm{x})$ on any input $\bm{x}\in\mathbb{R}^{n}$ arises solely through the transformed representation $\bm{W}_{1}^{\top}\bm{x}$.
@@
The technical contribution underlying Proposition 1 and Theorem 1 is to extend the orthogonal-invariance-based theory of marginal asymptotic normality—originally developed by zhao2022asymptotic and later shown to be broadly applicable by sawaya2024high—from loss minimizers (i.e., M-estimators) to the individual iterates of loss-minimization algorithms. As a consequence, our asymptotic normality results do not rely on
@@
The availability of early stopping is also appealing, as it may help reduce type-II error. Points (ii) and (iii) are particularly important in the training of deep neural networks, where convergence to a global minimum is rarely guaranteed. Furthermore, our proofs extend naturally to gradient-based procedures for fitting conventional M-estimators without neural networks. This implies that valid inference remains possible even in settings where a maximum likelihood estimator may fail to exist—as in logistic regression [candes2020phase]—which further underscores the practical value of our approach.
@@ 4. Methodology
According to Theorem 1, $\xi_{j}^{(t)}$ satisfies (a) but not (b), whereas $|\xi_{j}^{(t)}|$ satisfies (b) but not (a). Therefore, we adopt a data-splitting approach [dai2023false, dai2023scale]. Specifically, we randomly divide the data into two equal parts and, using the quantities $\xi_{j1}^{(t)}$ and $\xi_{j2}^{(t)}$ computed from each split, construct the importance statistics, for $j\in[n]$,
@@
Such assumptions frequently appear in the related literature. Compared with Assumption 3.2 in dai2023scale, which requires that a fixed proportion of the true signals diverge, our assumption can be regarded as considerably weaker.
@@
A drawback of the data-splitting method is that it effectively halves the sample size, which may reduce power. One remedy, known in the literature and also applicable here, is to aggregate the selection results obtained from multiple random splits of the data [dai2023false]. Related stabilization techniques include du2023false and ren2024derandomised
@@ 5.3. Comparison with other methods
We benchmark our approach against flexible variable-selection baselines that can (or aim to) control FDR. Neural baselines advertised as enabling FDR control via deep representations include the Neural Gaussian Mirror (NGM) [xin2020ngm] and DeepLINK [zhu2021deeplink]. While empirically competitive, these methods do not offer formal guarantees. All methods are run at the same nominal FDR level $\alpha=0.1$.
We ran DeepLINK [zhu2021deeplink] with authors’ official code [^1] https://github.com/zifanzhu/DeepLINK with hyperparameters as follows: L1 penalty $10^{-3}$, learning rate $10^{-3}$, ELU activation, and mean squared error loss. Column-wise centering and scaling were applied by default.
@@ 6. Discussion
In this work, we considered the raw input gradients of a trained neural network as a feature importance. A promising direction for future work is to extend our analysis to more sophisticated feature attribution methods, including Integrated Gradients [sundararajan2017axiomatic], DeepLIFT [shrikumar2017learning], and SmoothGrad [smilkov2017smoothgrad]. Adapting our FDR-control framework to such attribution methods would potentially yield more stable and interpretable feature selection.
@@ A.1. Proof of Proposition 1
The basic idea of the proof is inspired by the proof of Proposition 2.1 in zhao2022asymptotic. Since $\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}/\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|$ has the unit norm and lies in $\mathrm{Col}(\bm{B})^{\perp}$, it is sufficient to show that, for any orthogonal matrix $\bm{U}\in\mathbb{R}^{n\times n}$ obeying $\bm{U}\bm{B}=\bm{B}$,
@@ A.2. Proof of Theorem 1
Since $\mathsf{rank}(\bm{B})=q^{*}$ by Assumption 1 (i), we have $\|\bm{P}_{\bm{B}}^{\perp}\|_{F}^{2}=n-q^{*}$ and $\|\bm{P}_{\bm{B}}^{\perp}\|_{\rm op}=1$. Thus, Hanson–Wright inequality [vershynin2018high] implies that, for any $t>0$,
@@
$\mathbb{P}\left(z_{j}<\frac{u}{1+\epsilon}\right)\leq\mathbb{P}\left(\sigma_{n}z_{j}\mid E_{n}\right)\leq\mathbb{P}\left(z_{j}<\frac{u}{1-\epsilon}\right).$ (40)
@@ A.3. Proof of Thorem 2
Our argument is inspired by the proof of Proposition 3.2 of dai2023scale, but proceeds under weaker conditions on signal strength and dimensionality. While auxiliary lemmas overlap with dai2023scale, we provide full proofs to ensure a self-contained presentation. We prepare some notation for the proof.
@@
$=\mathbb{E}\left[Q(I_{u}(x))Q(I_{u}(y))\mid\varsigma_{j^{\prime}1}^{(t)}=x>0,\,\varsigma_{j1}^{(t)}=y>0\right]+2\Delta.$ (72)
Similarly, we can upper bound $I_{2}$, $I_{3}$, and $I_{4}$. Combining the four upper bounds together, we obtain an upper bound on $\mathbb{P}(\breve{M}_{j}>u,\breve{M}_{j^{\prime}}>u)$ as

### v2

# Provable FDR Control for Deep Feature Selection: Deep MLPs and Beyond

@@ Abstract.
We develop a flexible feature selection framework based on deep neural networks that approximately controls the false discovery rate (FDR), a measure of Type-I error. The method applies to architectures whose first layer is fully connected. From the second layer onward, it accommodates multilayer perceptrons (MLPs) of arbitrary width and depth, convolutional and recurrent networks, attention mechanisms, residual connections, and dropout. The procedure also accommodates stochastic gradient descent with data-independent initializations and learning rates. To the best of our knowledge, this is the first work to provide a theoretical guarantee of FDR control for feature selection within such a general deep learning setting.
@@ 1. Introduction
On the statistics side, a central objective has been to provide theoretical guarantees on false discovery rate (FDR) control [1, 15, 2, 6, 31, 9, 8, 10], often by debiasing sparse estimators like the LASSO [28]. These guarantees, albeit rigorous, typically come at the expense of strong modeling assumptions (e.g., the true model follows a generalized linear model, or the feature distribution is known), which can diverge from complex real‑world phenomena and thereby cast doubt on the reliability of the selected features.
On the XAI side, a wealth of attribution methods, e.g., LIME [19], SHAP [14], random forest feature importance [5], and saliency maps [24], quantify the contribution of individual features in black‑box models. Thresholding such scores yields a practical selection heuristic, yet without guarantees on Type‑I error; thus, error control has remained elusive.
@@ 1.1. Related works
There are two main research streams of FDR control in feature selection: estimator-based FDR control and knockoff filters. In the former, e.g., for linear regression, whether coefficients are zero or nonzero determines which features should be selected, and FDR control can be achieved by invoking asymptotic normality of the estimators. Representative examples include the Gaussian mirror [31] and data splitting [9]. A key advantage of these approaches is scale-free property, since they avoid estimating the asymptotic variance, and they have been extended to generalized linear models [8]. Similar techniques have also been applied to sliced inverse regression under multi-index models [33], although the theoretical assumptions on the feature dimension and the sparsity level are rather stringent.
As the second line of research, the model-X knockoff [6] is groundbreaking in that it can control the FDR without assuming the structure of $\bm{y}\mid\bm{X}$ where $\bm{y}\in\mathbb{R}^{m}$ is a response vector. However, it requires knowledge of the joint distribution of $\bm{X}$, which is restrictive. To address this limitation, methods that estimate the distribution of $\bm{X}$ using generative networks have been proposed [11, 20]. In addition, a sequential knockoff sampler for a given feature distribution has been proposed [3].
@@
Several studies have proposed FDR control methods that employ neural networks, although without theoretical guarantees and typically within restricted classes of architectures. The Neural Gaussian Mirror [30] defines a kernel-based conditional dependence measure and performs feature selection with MLPs. DeepPINK [13] is a knockoff framework with a specially designed network architecture in which the features are assumed to be jointly Gaussian, and DeepLINK [34] relaxes this distributional restriction. These approaches have been shown empirically to achieve FDR control. Nevertheless, because theoretical guarantees are not provided, it remains unclear under what conditions FDR control is achievable, and the range of supported network architectures is limited.
@@ 2. Problem formulation
Suppose that we observe a response vector $\bm{y}=(y_{1},\ldots,y_{m})^{\top}\in\mathbb{R}^{m}$ together with a design matrix $\bm{X}=(\bm{x}_{1},\ldots,\bm{x}_{m})^{\top}\in\mathbb{R}^{m\times n}$, where $m$ is the sample size and $n$ is the number of features. Our goal is to select a relevant subset of feature indices from $[n]$ that are associated with $\bm{y}$. A desirable feature selection procedure controls the false discovery rate (FDR) [4], defined by
@@ 3. Theoretical background
The first hidden layer of the network exists and is taken to be dense and fully connected, and we denote its weight matrix by $\bm{W}_{1}\in\mathbb{R}^{n\times q}$. The dependence of the entire network $f_{\mathcal{W}}(\bm{x})$ on any input $\bm{x}\in\mathbb{R}^{n}$ arises solely through the transformed representation $\bm{W}_{1}^{\top}\bm{x}$.
@@
The technical contribution underlying Proposition 1 and Theorem 1 is to extend the orthogonal-invariance-based theory of marginal asymptotic normality—originally developed by Zhao et al. [32] and later shown to be broadly applicable by Sawaya et al. [21]—from loss minimizers (i.e., M-estimators) to the individual iterates of loss-minimization algorithms. As a consequence, our asymptotic normality results do not rely on
@@
The availability of early stopping is also appealing, as it may help reduce type-II error. Points (ii) and (iii) are particularly important in the training of deep neural networks, where convergence to a global minimum is rarely guaranteed.
Example 3.1 (Linear regression).
Among neural networks satisfying Assumption 3, the simplest example is the linear model $f_{\mathcal{W}}(\bm{x})=\bm{w}_{1}^{\top}\bm{x}$, i.e., single-layer neural networks. Since the multi-index model in Definition 1 includes the linear regression model $y_{i}=\bm{\beta}^{\top}\bm{x}_{i}+\varepsilon_{i}$ satisfying $\mathbb{E}[\varepsilon_{i}\mid\bm{x}_{i}]=0$, our result in Theorem 1 can be viewed as extending asymptotic normality results in high dimensions for arbitrary $M$-estimators satisfying Assumption 2 in linear regression to every iteration of stochastic gradient descent.
Example 3.2 (Logistic regression).
The multi-index model in Definition 1 also encompasses the logistic regression model $y_{i}=g(\bm{\beta}^{\top}\bm{x}_{i},\varepsilon_{i})$, where $g(u,v)=\mathbbm{1}\{(1+e^{-u})^{-1}>v\}$ and $\varepsilon_{i}\overset{\mathrm{iid}}{\sim}\mathrm{Unif}(0,1)$ independent of $\bm{x}_{i}$. By taking $f_{\mathcal{W}}(\bm{x})=\bm{w}_{1}^{\top}\bm{x}$ and the cross-entropy loss $\mathcal{L}(u,v)=\log(1+\exp(v))-uv$, our assumptions also cover the gradient descent iterates used to compute the maximum likelihood estimator. As shown by Candès and Sur [7], the maximum likelihood estimator in logistic regression may fail to exist. Notably, because our theory applies to each iterate of SGD, it remains valid even in such settings where the optimization trajectory diverges and convergence is not guaranteed.
@@ 4. Methodology
According to Theorem 1, $\xi_{j}^{(t)}$ satisfies (a) but not (b), whereas $|\xi_{j}^{(t)}|$ satisfies (b) but not (a). Therefore, we adopt a data-splitting approach [9, 8]. Specifically, we randomly divide the data into two equal parts and, using the quantities $\xi_{j1}^{(t)}$ and $\xi_{j2}^{(t)}$ computed from each split, construct the importance statistics, for $j\in[n]$,
@@
Such assumptions frequently appear in the related literature. Compared with Assumption 3.2 in Dai et al. [8], which requires that a fixed proportion of the true signals diverge, our assumption can be regarded as considerably weaker.
@@
A drawback of the data-splitting method is that it effectively halves the sample size, which may reduce power. One remedy, known in the literature and also applicable here, is to aggregate the selection results obtained from multiple random splits of the data [9], as is commonly done in stability selection [16, 22]. Related stabilization techniques include Du et al. [10] and Ren and Barber [18]. Later, Takahashi [27] theoretically shows that such stabilization techniques increase power in the proportional asymptotics.
@@ 5.3. Comparison with other methods
We benchmark our approach against flexible variable-selection baselines that can (or aim to) control FDR. Neural baselines advertised as enabling FDR control via deep representations include the Neural Gaussian Mirror (NGM) [30] and DeepLINK [34]. While empirically competitive, these methods do not offer formal guarantees. All methods are run at the same nominal FDR level $\alpha=0.1$.
We ran DeepLINK [34] with authors’ official code [^1] https://github.com/zifanzhu/DeepLINK with hyperparameters as follows: L1 penalty $10^{-3}$, learning rate $10^{-3}$, ELU activation, and mean squared error loss. Column-wise centering and scaling were applied by default.
@@ 6. Discussion
In this work, we considered the raw input gradients of a trained neural network as a feature importance. A promising direction for future work is to extend our analysis to more sophisticated feature attribution methods, including Integrated Gradients [26], DeepLIFT [23], and SmoothGrad [25]. Adapting our FDR-control framework to such attribution methods would potentially yield more stable and interpretable feature selection.
@@ A.1. Proof of Proposition 1
The basic idea of the proof is inspired by the proof of Proposition 2.1 in Zhao et al. [32]. Since $\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}/\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|$ has the unit norm and lies in $\mathrm{Col}(\bm{B})^{\perp}$, it is sufficient to show that, for any orthogonal matrix $\bm{U}\in\mathbb{R}^{n\times n}$ obeying $\bm{U}\bm{B}=\bm{B}$,
@@ A.2. Proof of Theorem 1
Since $\mathsf{rank}(\bm{B})=q^{*}$ by Assumption 1 (i), we have $\|\bm{P}_{\bm{B}}^{\perp}\|_{F}^{2}=n-q^{*}$ and $\|\bm{P}_{\bm{B}}^{\perp}\|_{\rm op}=1$. Thus, Hanson–Wright inequality [29] implies that, for any $t>0$,
@@
$\mathbb{P}\left(z_{j}<\frac{u}{1+\epsilon}\right)\leq\mathbb{P}\left(\sigma_{n}z_{j}<u\mid E_{n}\right)\leq\mathbb{P}\left(z_{j}<\frac{u}{1-\epsilon}\right).$ (40)
@@ A.3. Proof of Thorem 2
Our argument is inspired by the proof of Proposition 3.2 of Dai et al. [8], but proceeds under weaker conditions on signal strength and dimensionality. While auxiliary lemmas overlap with Dai et al. [8], we provide full proofs to ensure a self-contained presentation. We prepare some notation for the proof.
@@
Under the assumptions of Theorem 1, as $n\to\infty$ while $q^{*}=o(n)$, we have,
$\sup_{i,j\in S^{\mathsf{c}},\,t_{1},t_{2}\in\mathbb{R}}\left\lvert\mathbb{P}\left(\frac{\sqrt{n}\xi_{i}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}<t_{1},\frac{\sqrt{n}\xi_{j}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}<t_{2}\right)-\Phi(t_{1})\Phi(t_{2})\right\rvert\to 0.$ (58)
Proof of Lemma A.5.
We assume, without loss of generality, that $t_{1}>0$ and $t_{2}>0$. Using the fact that ${\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}\overset{\rm d}{=}{\bm{P}_{\bm{B}}^{\perp}\bm{z}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{z}\|}$ where $\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$ by Theorem 1, we have,
$\left(\frac{\sqrt{n}\xi_{i}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|},\frac{\sqrt{n}\xi_{j}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}\right)\overset{\rm d}{=}\left(\frac{\sqrt{n}z_{i}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{z}\|},\frac{\sqrt{n}z_{j}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{z}\|}\right),$ (59)
for $i,j\in S^{\mathsf{c}}$. This follows from the fact that the $j$-th element of ${\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}$ is ${\xi_{j}^{(t)}}/{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}$ under the null since $\bm{b}_{j}=\bm{0}$ for $j\in S^{\mathsf{c}}$. Denote $\sigma_{n}=\sqrt{n}/\|\bm{P}_{\bm{B}}^{\perp}\bm{z}\|$. From (59), it follows that
$\Delta_{n}$ $\equiv\sup_{i,j\in S^{\mathsf{c}},\,t_{1},t_{2}\in\mathbb{R}}\left\lvert\mathbb{P}\left(\frac{\sqrt{n}\xi_{i}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}<t_{1},\frac{\sqrt{n}\xi_{j}^{(t)}}{\|\bm{P}_{\bm{B}}^{\perp}\bm{\xi}^{(t)}\|}<t_{2}\right)-\Phi(t_{1})\Phi(t_{2})\right\rvert$ (60)
$=\sup_{i,j\in S^{\mathsf{c}},\,t_{1},t_{2}\in\mathbb{R}}\left\lvert\mathbb{P}\left(z_{i}<\sigma_{n}t_{1},z_{j}<\sigma_{n}t_{2}\right)-\Phi(t_{1})\Phi(t_{2})\right\rvert$. (61)
Fix an arbitrary $\epsilon>0$ and denote $E_{n}=\{|\sigma_{n}-1|<\epsilon\}$. Then we have
$\left\lvert\mathbb{P}\left(z_{i}<\sigma_{n}t_{1},z_{j}<\sigma_{n}t_{2}\right)-\Phi(t_{1})\Phi(t_{2})\right\rvert$ (62)
$\leq\mathbb{P}(E_{n}^{\mathsf{c}})+\left\lvert\mathbb{P}\left(z_{i}<\sigma_{n}t_{1},z_{j}<\sigma_{n}t_{2}\mid E_{n}\right)-\Phi(t_{1})\Phi(t_{2})\right\rvert$. (63)
$\mathbb{P}\left(z_{i}<\frac{t_{1}}{1+\epsilon},z_{j}<\frac{t_{2}}{1+\epsilon}\right)\leq\mathbb{P}\left(z_{i}<\sigma_{n}t_{1},z_{j}<\sigma_{n}t_{2}\mid E_{n}\right)\leq\mathbb{P}\left(z_{i}<\frac{t_{1}}{1-\epsilon},z_{j}<\frac{t_{2}}{1-\epsilon}\right).$ (64)
From this, we have
$\left\lvert\mathbb{P}\left(z_{i}<\sigma_{n}t_{1},z_{j}<\sigma_{n}t_{2}\mid E_{n}\right)-\Phi(t_{1})\Phi(t_{2})\right\rvert$ (65)
$\leq\max\left\{\Phi\left(\frac{t_{1}}{1-\epsilon}\right)\Phi\left(\frac{t_{2}}{1-\epsilon}\right)-\Phi(t_{1})\Phi(t_{2}),\,\Phi(t_{1})\Phi(t_{2})-\Phi\left(\frac{t_{1}}{1+\epsilon}\right)\Phi\left(\frac{t_{2}}{1+\epsilon}\right)\right\}.$ (66)
Taking the supremum, the mean-value theorem gives
$\sup_{t_{1},t_{2}\in\mathbb{R}}\left\lvert\mathbb{P}\left(z_{i}<\sigma_{n}t_{1},z_{j}<\sigma_{n}t_{2}\mid E_{n}\right)-\Phi(t_{1})\Phi(t_{2})\right\rvert\leq\frac{\epsilon}{1-\epsilon}\cdot\sqrt{\frac{2}{\pi e}},$ (67)
$\Delta_{n}\leq\mathbb{P}(|\sigma_{n}-1|>\epsilon)+\frac{\epsilon}{1-\epsilon}\cdot\sqrt{\frac{2}{\pi e}}.$ (68)
Lemma A.6.
@@
$=\mathbb{E}\left[Q(I_{u}(x))Q(I_{u}(y))\mid\varsigma_{j^{\prime}1}^{(t)}=x>0,\,\varsigma_{j1}^{(t)}=y>0\right]+2\Delta,$ (83)
where the inequality follows from Lemma A.5. Similarly, we can upper bound $I_{2}$, $I_{3}$, and $I_{4}$. Combining the four upper bounds together, we obtain an upper bound on $\mathbb{P}(\breve{M}_{j}>u,\breve{M}_{j^{\prime}}>u)$ as